OpenStudy (lgbasallote):

An athlete swims the length \(L\) of a pool in a time \(t_1\) and makes the return trip to the starting position in a time \(t_2\). If she is swimming initially in the positive x direction, determine her average velocities symbolically in (a) the first half of the swim help me analyze this @rajathsbhat ? :)

OpenStudy (anonymous):

OK

OpenStudy (lgbasallote):

do we use speed = distance/time?

OpenStudy (anonymous):

yes, but that's not the ques.

OpenStudy (lgbasallote):

aww :9

OpenStudy (lgbasallote):

:( *

OpenStudy (anonymous):

It's just that the avg velocity of the swimmer is positive in the first half & -ve in the second. That's all.

OpenStudy (lgbasallote):

i see i see..makes sense...so how?

OpenStudy (anonymous):

You see they have asked you to take the direction in which she starts to swim as the +ve X direction

OpenStudy (anonymous):

so deltax will be +ve during that time.

OpenStudy (lgbasallote):

uhh maybe the solution can make sense? o.O

OpenStudy (lgbasallote):

im not really good with terms...im better in analyzing algebra lol

OpenStudy (anonymous):

I'll try to draw you a diagram

OpenStudy (lgbasallote):

ugh please no diagrams o.O just solution

OpenStudy (lgbasallote):

diagrams tend to confuse me a lot...the only time i was able to get it was when i studied calculus and it took me a hard time

OpenStudy (anonymous):

I'll write it with my own hand & scan it

OpenStudy (lgbasallote):

uhmm sure...ill try to understand it...but i really think the solution can help me better o.O

OpenStudy (anonymous):

Well, as for thee sol, it is "positive"

OpenStudy (lgbasallote):

do we use distance/time?

OpenStudy (lgbasallote):

is it \[\frac{L}{t_1}\]

OpenStudy (anonymous):

No, we use v=delta x/t1 =+L/t1

OpenStudy (lgbasallote):

so just add +?

OpenStudy (anonymous):

For the return swim, it would be v=-L/t1

OpenStudy (anonymous):

Yes, cause the direction is very important in this Ques

OpenStudy (lgbasallote):

shouldnt it be \[-\frac{L}{t_2}?\]

OpenStudy (anonymous):

Ahh, sorry, yes, it should.

mathslover (mathslover):

average velocity for the first half of the swim= L/T1

OpenStudy (lgbasallote):

what about for round trip? \[\frac{2L}{t_2 - t_1}?\]

mathslover (mathslover):

0 by velocity she did not change position

OpenStudy (anonymous):

No, the avg vel for the round trip is 0

OpenStudy (lgbasallote):

what about average speed?

OpenStudy (lgbasallote):

that's the correct symbolism?

mathslover (mathslover):

average speed for the round trip= 2L/(T1+T2)

OpenStudy (lgbasallote):

oh..plus

OpenStudy (anonymous):

2L/t1+t2

mathslover (mathslover):

yes u missed ... no problem

OpenStudy (lgbasallote):

great//thanks ^_^

OpenStudy (anonymous):

Awesome.

OpenStudy (anonymous):

yw :)

mathslover (mathslover):

well as u commented there diagrams make the question complicated for you .. i will prefer you to just draw a simple diagram first ... like that of this one

mathslover (mathslover):

in physics without diagrams there is nothing to analyze with. :(

OpenStudy (anonymous):

true, mathslover, very true.